OpenStudy (anonymous):

Prove (x+y)(x+y)=x^2+2xy+y^2. Could you please check our work? I'm working with a partner on this and I have to be sure we are correct. Thank you.

OpenStudy (anonymous):

x/y + y/x = 2 Given. x≠0 and y≠0 Because then the original question would be dividing by zero. xy≠0 Because neither factor is zero. (xy) (x/y + y/x) = (xy) 2 Multiply both sides of given equation by (xy); we can do this because we are not multiplying by zero. x2 + y2 = 2xy Simplify. x2 - 2xy + y2 = 0 Make one side zero. (x-y)2 = 0 Factor it. (x-y) = 0 Take the square root of both sides. x = y. Solved.

jimthompson5910 (jim_thompson5910):

for your first part, all I see are boxes instead of variables (see attached)

OpenStudy (anonymous):

Thank you for telling me. Just one sec.

OpenStudy (anonymous):

you see now?

jimthompson5910 (jim_thompson5910):

ok much better

OpenStudy (anonymous):

I also think I could use FOIL with the left side to make it equal the right side. That seems the simplest to me.

jimthompson5910 (jim_thompson5910):

yes you can use FOIL or the distributive property go from (x+y)(x+y) to x^2+2xy+y^2

OpenStudy (alexandervonhumboldt2):

easy: (x+y)*(x+y)=x^2+xy+yx+^2=x^2+2xy+y^2

OpenStudy (alexandervonhumboldt2):

simplify FOIL it

jimthompson5910 (jim_thompson5910):

you can also use the box method as a visual way to do it

jimthompson5910 (jim_thompson5910):

To be honest, I'm not sure how the two parts are connected. I don't see (x+y)(x+y) anywhere in the second part. I guess it's similar to (x-y)^2 = (x-y)(x-y) = x^2-2xy+y^2

OpenStudy (anonymous):

Could you show me the box method please?

jimthompson5910 (jim_thompson5910):

sure we have (x+y) times (x+y) there are 2 terms in each parenthesis so we will have a 2x2 box like this 

jimthompson5910 (jim_thompson5910):

along the left and top edges we place the terms from x+y and x+y